Suppose that Z is a continuous random variable. Let b(2) denote the unnormalized PDF of Z ―the function \phi satisfies all properties of a PDF except that it is not normalized. Now suppose we use \phi to compute something like the moment generating function (MGF), i.e., we compute the function

\gamma (s)= \int_{-\infty}^{\infty} e^{sz}\phi(z)dz

What is \gamma (0)? How can we use \gamma (0) to normalize the PDF?
Answer #1

let constant = k

such that

\int_{-\infty }^{\infty }k \phi(z)dz = 1

\Rightarrow k \int_{-\infty }^{\infty } \phi(z)dz = 1

h(0) = 1

k= γ(0)

hence

we have divide b(2) by \gamma (0)   to normalize the pdf